OpenStudy (australopithecus):

An object moves along the x axis according to the equation x(t) = (3.00t2 – 2.00t + 3.00) m, where t is in seconds. Determine (a) the average speed between t = 2.00 s and t = 3.00 s, (b) the instantaneous speed at t = 2.00 s and at t = 3.00 s, (c) the average acceleration between t = 2.00 s and t = 3.00 s, and (d) the instantaneous acceleration at t = 2.00 s and t = 3.00 s. so for a) I got x(2) = 11 x(3) = 24 b) x'(2) = 10m/s x'(3) = 16m/s c) Vxavg = Delta x/ Delta t = (24-11)/(3-2) = 13m/s d) x''(t) = 6 I'm probably wrong for most of these answers can anyon

OpenStudy (australopithecus):

*anyone clarify it cut me off

OpenStudy (anonymous):

a) I got x(2) = 11 x(3) = 24 Vxavg = Delta x/ Delta t = (24-11)/(3-2) = 13m/s

OpenStudy (anonymous):

b) x'(2) = 10m/s x'(3) = 16m/s

OpenStudy (anonymous):

c and d) x''(t) = 6 because a is constant for all t a=6 m/s^2

OpenStudy (australopithecus):

I though velocity was displacement/time

OpenStudy (australopithecus):

so wouldn't it be 24-11/3-2

OpenStudy (anonymous):

yes thats right

OpenStudy (australopithecus):

and d) is the second derivative of the function of position so d) is 6

OpenStudy (anonymous):

thats right

OpenStudy (australopithecus):

alright thanks for confirming for me

OpenStudy (anonymous):

welcome dear